# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

### v9

@@ 2.2. Multi-leaf Complex Manifolds of Algebroid Functions
for Similarly, we obtain
for Hence, is a biholomorphic mapping. On the other hand, for any there corresponds to distinct function elements Therefore, we conclude that
@@
In what follows, we shall construct a -leaf complex manifold determined by all functions elements subordinate to equation (7). Let be any two function elements subordinate to equation (7). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branching set of for Set
Remark 2.18.
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in where is a -order branching point of satisfying In other words, can be separated into single-valued components for
@@
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one can regard that
@@
all such that is a direct analytic continuation of satisfying
all satisfying
where is a neighborhood of In particular, a -neighborhood of (satisfying ), is defined by the set which contains the following elements:
all such that is a direct analytic continuation of satisfying
all satisfying
A subset is called an open set, if either is an empty set or every point is an inner point, namely, there exists a -neighborhood of such that If is a -leaf function element, then has leaves. In this case, we call a -leaf neighborhood of This gives a topology of and this topology contains the topology of
@@
Hausdorff property of can be showed using the similar arguments as in the proof of Theorem 4.3. Next, we prove that is connected. If not, one can write where are two disjoint open subsets of Set It is evident that Notice that are two disjoint open subsets of since is an analytic set of complex codimension This means that is not connected, but it is a contradiction. ∎
Next, let us continue to endow with a complex structure such that is a complex manifold. Take any -leaf function element Assume that is a non-singular branching point of Apply Theorem 2.14 ( is properly chosen if necessary), there exists a biholomorphic mapping such that can be expanded into the Pusieux series
satisfying that and for
@@
all such that is a direct analytic continuation of satisfying
all satisfying
Let be single-valued connected components of Define a mapping by
where Note that is a homeomorphism and its inverse is
where is a positive integer such that Let be the set of all function elements such that is a singular branching point of some multi-valued component of and let be the set of all corresponding singular branching points. It is evident that is an analytic set of complex dimension not greater than Set
We will show that is a complex atlas of Suppose It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf meromorphic function element and is a regular function element. For every there exists a single-valued component of such that in some neighborhood of i.e., It is clear that because all function elements in are meromorphic. A direct computation leads to
where is a positive integer such that Hence, is a holomorphic mapping. On the other hand, its inverse is
@@
Finally, we treat the singular part Let be the natural projection by mapping a function element into Then, we note that is a generic -sheeted analytic covering such that
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via To conclude, we obtain:
@@
is a non-compact -leaf complex manifold of complex dimension which is called the complex manifold of
@@ 2.3. Zero and Pole Divisors of Algebroid Functions
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors. Thus, we assume that It yields that
@@
Combining this with (10), we see that reaches the maximal order of a pole of Since then we obtain
which follows that ∎
@@
Proof.
Set Apply the similar arguments as in the proof of Theorem 2.22, we can prove the theorem. ∎
@@ 3. Uniformization of Algebroid Functions
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold. Our main idea comes from the excellent work of W. Stoll [35] on algebroid reduction.
@@ 3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, namely, it is a proper and surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex dimension It is evident that
is a proper, surjective, local biholomorphic mapping of sheet number and hence a -sheeted analytic covering. By Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let the mapping degree of at is defined in such manner: if is a relatively compact neighborhood of such that then
@@
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection mapping. Let be the -fold symmetric tensor product of i.e.,
@@
for any is called the symmetric tensor product of and is the -order symmetric group. It is evident that is a subspace of -fold tensor product space with Again, put
@@
Let be a meromorphic mapping with indeterminacy set Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
@@ 3.2. Uniformization
Let be the maximal integer in such that Then, equation (13) reduces to Let be the roots of the equation. Apply Vieta’s theorem, it is immediate that
@@
By this with (11), we have To conclude, we obtain
@@
Let be defined by equation (7). It gives a differentiably non-degenerate and irreducible meromorphic mapping
@@
By
we have the theorem proved. ∎
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of generic sheet number via which is a proper surjective holomorphic mapping.
@@
Let be the -leaf complex manifold of a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
@@ 4. Nevanlinna Theory of Meromorphic Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
in local holomorphic coordinates Moreover, we define the Ricci form of by
@@ 4.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point and denote by the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
@@
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
@@
Apply Dynkin formula (see Lemma 4.1) to we obtain
@@ 4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic with non-negative Ricci Curvature. Fix a reference point and denote by the geodesic ball centered at with radius in Denote by the Riemannian volume of The non-parabolicity of means that
@@ 4.2.2. Gradient estimates of Green functions
Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [8] proved the following theorem.
@@
By
we obtain an estimate for upper bounds of as follows.
@@ 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
@@
in local holomorphic coordinates here is the inverse of Define the Nevanlinna’s characteristic function of by
@@
Whence, it concludes from Fubini’s theorem that
@@ 4.2.4. Second Main Theorem
Define a non-negative function on via
@@
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1)
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and
@@
For any we can take local holomorphic coordinates near and a local holomorphic coordinate near such that
@@
holds for all outside where one used the fact that Combining (4.2.4) and (19), we have the theorem proved. ∎
@@ A. Setups for
Due to the curvature conditions, we have a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
@@
Since as and then we deduce that is a precompact domain containing in which exhausts So, the boundary of can be formulated as
@@ B. Setups for
Notice that is parabolic when i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case. However, we note (21) for it is thus natural to define
for Let it then yields from (22) that
@@
Let be a meromorphic function. Similarly, one can defined the Nevanlinna’s functions and of as well as on as in Section 4.2.1 and Section 4.2.4.
@@ 4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
When is the geodesic ball centered at with radius The conclusion holds immediately due to (23). In the following, we consider the other case. Since on then for
@@
When is a geodesic ball centered at with radius Hence, it yields from (4) and (23) that
@@ 5. Nevanlinna Theory of Algebroid Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
@@ 5.1. First Main Theorem
Let be an arbitrary relatively compact domain with smooth boundary in Fix a point then we have Green function for and harmonic measure on (see definitions in Section 4.1).
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
Here, are single-valued meromorphic components of namely, are locally meromorphic on outside Recall that it defines a meromorphic mapping with indeterminacy locus We can choose such that
@@
With the aid of Vieta’s theorem, we obtain Dynkin formula yields that
By definition, the left hand side equals
@@
Set then
with
In particular, and Similarly, we have
which means that
Since
then we obtain
@@
5.2. Estimates of Branching Divisors
Let be a -valued algebroid function on We use the same notations such as given in Section 5.1. Our aim is to give an estimate of the branching divisor of Denote by the -valued, -valued components of at any respectively. Then, we have namely, is a -order branching point of for and is hence a -order branching point of The branching divisor of associated to is defined by a divisor on satisfying that
@@ 5.2. Estimates of Branching Divisors
Review that the discriminant of is given by
@@
Take any Let be the -valued, -valued components of at respectively, where Since the set of singular branching points of is an analytic set of complex dimension not greater than then one just needs to consider that is a non-singular branching point of By Theorem 2.14, we can write for
in a local holomorphic coordinate near such that has a local coordinate If then is not a pole of So, one can derive from an observation over that there exist terms in which contain the factor i.e., the factor appears at least times in It implies that If then is a pole of By Theorem 2.22, we have
This implies the lemma holds. ∎
@@
Theorem 5.3 (Estimates of branching divisors).
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold of a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.21. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula (see Lemma 4.1) still holds under the pullback metric
@@
Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then we may assume that (otherwise, a subsequence can be chosen so that as ). Let be the Laplace-Beltrami operators of on respectively. We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively. It is not difficult to show that as and (weakly) uniformally on as due to
for any Borel subset of Similarly, we also have that (weakly) uniformally on as where denote the volume measures of for respectively. To unify the form, one uses to stand for the volume elements of for respectively. Apply Dynkin formula, it gives
@@
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex dimension
@@
Then, it is trivial to see that is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and also is the harmonic measure on with respect to Set
@@
Note that is a -sheeted analytic covering, and two connected components () might meet only at the branching set of and a branching point of is possibly a point in On the other hand, if is a branching point with order of we have multi-leaf function elements subordinate to equation (7), say where with and is a branching point of with order i.e., So, it is not hard to deduce from the construction of and the definition of that
On the other hand, we have and
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
holds for all outside Hence, we have the theorem proved. ∎
When we have By it is trivial to deduce Hence, we obtain:
@@
Assume that has non-positive sectional curvature satisfying (3). Note that the curvatures of also satisfy (3) under the pullback metric In the same way, we obtain notations and by setting Apply the similar arguments as in the proof of Theorem 5.4, we can show that
@@
If then and Hence, we conclude that
@@
in holomorphic global coordinates where
It is known that is a non-compact complete Kähler manifold of constant sectional curvature under this metric. From (29), the associated Laplace-Beltrami operator is
@@
where
Let be the Riemannian area element of Using the second formula in (31) and volume form formula, it gives immediately
@@
With these preparations, we shall give an estimate for Ricci term Let be the scalar curvature of which is a negative constant. It follows from (32) that
Substituting (30) into the expression of and calculating the integral, it is not hard to deduce that

### v10

@@ 2.2. Multi-leaf Complex Manifolds by Algebroid Functions
for any Similarly, we have
It shows that is a biholomorphic mapping. On the other hand, each corresponds to distinct function elements That is, we conclude that
@@
In what follows, we would construct a -leaf complex manifold determined by all function elements subordinate to equation (7). Let be any two function elements subordinate to equation (7). We define an equivalent relation as follows: if and only if
for some two single-valued components of respectively, where is the branching set of with Set
@@
A meromorphic function element is called a direct analytic continuation of a multi-leaf function element if both and in a neighborhood of for some single-valued component of where is the branching set of If so, one may regard that
@@
all with satisfying that is a direct analytic continuation of
all with Here and before, is an arbitrary neighborhood of
In particular, a -neighborhood of with is defined by the set that contains the following elements:
all with satisfying that is a direct analytic continuation of
all with
A subset is called an open set, if is an empty set or every point is an inner point, i.e., there exists a -neighborhood of such that It gives a topology of which contains the topology of If is a -leaf function element, then has leaves at We call a -leaf neighborhood of
@@
The Hausdorff property of is showed using the similar arguments as in the proof of Theorem 2.7. We assert that is connected. Otherwise, one may assume that where are two disjoint open subsets of Set It is evident that Since is an analytic set of complex codimension we see that are two disjoint open subsets of This implies that is not connected, which is a contradiction. ∎
Next, we will equip with a complex structure such that is a complex manifold. Taking a -leaf function element Assume that is not a singular branching point of and is chosen suitably. Thanks to Theorem 2.14, there exists a biholomorphic mapping such that can be expanded into the Pusieux series
on with and for all
@@
all with satisfying that is a direct analytic continuation of
all with
Let be distinct single-valued components of Define a mapping by
for , where We note that is a homeomorphism and its inverse is
where is an integer such that
Let be the set of all function elements satisfying that is a singular branching point of Moreover, denotes the set of all the corresponding singular branching points. It is clear that is an analytic set of complex codimension not smaller than Set
We now show that is a complex atlas of Assume that It is sufficient to verify that the transition mapping
is a biholomorphic mapping. It is automatically a homeomorphism. Through a simple analysis, we only need to deal with the situation in which is a -leaf meromorphic function element and is a regular function element. For any there exists a single-valued component of satisfying that in some neighborhood of i.e., It is clear that since all function elements in are meromorphic. A direct computation leads to
where is an integer such that
Thus, is a holomorphic mapping. On the other hand, its inverse is
@@
Treat the singular part Let be the natural projection by mapping any function element to Note that is a generic -sheeted analytic covering such that
Since is an analytic set of complex codimension not smaller than the complex structure of can be viewed as a natural extension of the complex structure of This implies that one can naturally extend the complex structure of to via Hence, we conclude that
@@
is a non-compact -leaf complex manifold of complex dimension We call the complex manifold by
@@ 2.3. Zero and Pole Divisors of Algebroid Functions
Let be a zero of Since in the sense of divisors, we may assume that It follows that
@@
Combining this with (11), we see that reaches the maximal order of a pole of Since we obtain
which means that ∎
Similarly, we can show by setting that
@@ 3. Uniformization of Algebroid Functions
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold. The main idea originates from W. Stoll [35].
@@ 3.1. Algebroid Reduction
Let be a non-compact complex manifold of complex dimension Let be a generic -sheeted analytic covering, by which one means a proper, surjective holomorphic mapping with branching set i.e., the set of all points such that is not locally biholomorphic at Note that is an empty set or an analytic set of pure complex codimension It is clear that
is a proper, surjective, local biholomorphic mapping of sheet number and thus a -sheeted analytic covering. By R. Remmert [27] (see Andreotti-Stoll [2] also), the exceptional set
is an analytic set satisfying that and Moreover, and for Let The mapping degree of at is defined in such manner: if is a neighborhood of such that then
@@
Let be a complex vector space of complex dimension with a basis Let be the projection of where is the natural projection. Let be the -fold symmetric tensor product of i.e.,
@@
is called the symmetric tensor product of and stands for the -order symmetric group. Note that is a subspace of -fold tensor product space with Again, put
@@
Let be a meromorphic mapping with indeterminacy locus Then, the algebroid reduction of via is a meromorphic mapping with indeterminacy locus defined by
@@ 3.2. Uniformization
Let be the maximal integer in such that Then, equation (14) reduces to Let be the roots of the equation. Applying Vieta’s theorem, it is immediate that
@@
By this with (12), we have Hence, we obtain
@@
Let be defined by equation (7). It gives a differentiably non-degenerate, irreducible meromorphic mapping
@@
Combining with then we prove the theorem. ∎
In conclusion, given a -leaf complex manifold by a -valued algebroid function on there exists the natural projection of generic sheet number given by which is a proper surjective holomorphic mapping. That is, we have shown that
@@
Let be the -leaf complex manifold by a -valued algebroid function on Then, can lift to a meromorphic function via the natural projection such that
@@ 4. Nevanlinna Theory of Meromorphic Functions
Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
in a local holomorphic coordinate Moreover, the Ricci form of is defined by
@@ 4.1. First Main Theorem
Let be any precompact domain with smooth boundary in Fix a point and denote by the Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
@@
Next, we introduce Nevanlinna’s functions. Let be a meromorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
@@
Applying Dynkin formula to we obtain
@@ 4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic with non-negative Ricci Curvature. Fix a reference point and denote by the geodesic ball centered at with radius in Let denote the Riemannian volume of The non-parabolicity of means that
@@ 4.2.2. Gradient estimates of Green functions
Let stand for the gradient operator on a Riemannian manifold. Cheng-Yau [8] proved the following theorem.
@@
Applying
we obtain an upper bound estimate of as follows.
@@ 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
@@
in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
@@
It follows from Fubini’s theorem that
By this with the First Main Theorem, we conclude that
@@ 4.2.4. Second Main Theorem
Define a non-negative function on by
@@
in the sense of currents, in which is the zero divisor of and is the ramification divisor of Using Dynkin formula, we have
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on and on as well as
@@
For any we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
@@
holds for all outside where one used the fact that Combining (4.2.4) with (20), we have the theorem proved. ∎
@@ A. Setups for
By the curvature conditions, there exists a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
@@
Since as and we deduce that is a precompact domain containing in which exhausts Thus, the boundary of can be formulated as
@@ B. Setups for
It is noted that is parabolic if i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case. However, we note (22) for it is thus natural to define
for Let then it yields from (23) that
@@ 4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
When is the geodesic ball centered at with radius The conclusion holds immediately due to (4.3.1). In the following, we consider the other case. Write
Since on we deduce that
Hence, for
@@
When is a geodesic ball centered at with radius Hence, it yields from (4) and (4.3.1) that
@@ 5. Nevanlinna Theory of Algebroid Functions
Let be a complete non-compact Kähler manifold of with Laplace-Beltrami operator associated to the Kähler metric We utilize the same notations such as etc. defined in Section 4.
@@ 5.1. First Main Theorem
Let denote a precompact domain with smooth boundary in Fix a point As defined in Section 4.1, we have the Green function for and the harmonic measure on
In what follows, we introduce Nevanlinna’s functions of algbroid functions. We are given a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
such that defines a meromorphic mapping with indeterminacy locus As we note that are the distinct single-valued meromorphic components of on outside Assume that
@@
According to the Vieta’s theorem, we obtain Combining Dynkin formula, it yields that
By the definition, the left hand side equals
@@
Then, we arrive at
Setting to get
where
In particular, and Similarly, we obtain
which leads to
Since
@@
5.2. Estimate of Branching Divisor
Let be a -valued algebroid function on We use the same notations as defined in Section 5.1. Our purpose is to obtain an estimate of the branching divisor of For each we use to denote the all -valued, -valued components of at respectively. As seen before, we have Namely, has branching order at with and has branching order at The branching divisor of associated to is defined to be the divisor on satisfying that
@@ 5.2. Estimate of Branching Divisor
Recall that the discriminant of is given by
@@
For any we use to denote the all distinct function elements of at with leaves respectively. It is noted that From the assumption, we deduce that Since the set of all singular branching points of has complex codimension not smaller than it is sufficient to treat the case where is a non-singular branching point of By Theorem 2.14, we may write for
in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case that for all we deduce that the factor appears a total of at least times in Hence,
If then is a pole of For the worst case where for all it yields from Theorem 2.21 that
Combining the above two cases, we prove the lemma. ∎
@@
Theorem 5.3 (Estimate of branching divisor).
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold by a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.20. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula still holds under the pullback metric
@@
Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then one may assume that (otherwise, a subsequence can be chosen so that as ). Let be the Laplace-Beltrami operators of on respectively. We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively. It is not difficult to show that as and (weakly) uniformally on as due to
as for a Borel subset of Similarly, we have (weakly) uniformally on as in which stand for the volume measures of respectively. Set and Using Dynkin formula, we are led to
@@
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex codimension
@@
Then, it is easy to infer that is the Green function of for with a pole at satisfying Dirichlet boundary condition, as well as that is the harmonic measure on with respect to Set
@@
Note that is a -sheeted analytic covering, and two distinct connected components () may cross only at those points in the branching set of If is a -order branching point of then there are multi-leaf function elements with subordinate to equation (7). As we know that Since a point in may be a branching point of it yields from the definition of that
On the other hand, we have as well as
Combining the above with (5.3.1), it follows that for any there exists a subset of finite Lebesgue measure such that
holds for all outside By Theorem 5.3, we prove the theorem. ∎
When we have By it is trivial to deduce that Hence, we obtain:
@@
Assume that has non-positive sectional curvature satisfying (3). Note that the curvature of also satisfies (3) under the pullback metric In the similar way, we obtain notations and by setting Applying the same arguments as in the proof of Theorem 5.4, we can show that
@@
When we deduce that and Thus, we conclude that
@@
in a global holomorphic coordinate where
It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. From (31), the associated Laplace-Beltrami operator is
@@
with Let be the Riemannian area element of By the second formula in (33) and volume form formula, it gives immediately
@@
With these preparations, we shall give an upper estimate of Let be the scalar curvature of which is a negative constant. It yields from (34) that
Substituting (32) into the expression of and computing the integral, it is not hard to deduce that